# Model-Free Generic Robust Control for Servo-Driven Actuation Mechanisms with Layered Insight into Energy Conversions

### v3

@@ V-A Generic Servo-Driven Actuation System Model
where $\alpha_{1}$,…,$\alpha_{4}$ are unknown positive coefficients. $F_{1}$,…, $F_{4}$ are unknown state-variant modeling and uncertainty terms. $D_{1}$,…,$D_{4}$ are time-variant external disturbances and load effects. By defining $\upsilon=0,1,2,3$, $Sat_{\upsilon}(u_{\upsilon})=s_{1\upsilon}u_{\upsilon}(t)+s_{2\upsilon}$ is a constraining function of the control input $u_{\upsilon}$, which was defined as [24]:
$&s_{1\upsilon}=\begin{cases}\frac{1}{|u_{\upsilon}(t)|+1},&\text{ if }u_{\upsilon}(t)\geq u_{max}\text{ or }u_{\upsilon}(t)\leq u_{min}\\ 1&\text{ if }u_{min}\leq u_{\upsilon}(t)\leq u_{max}\end{cases}\\ &s_{2\upsilon}=\begin{cases}u_{min}-\frac{u_{\upsilon}(t)}{|u_{\upsilon}(t)|+1},&\text{ if }u_{\upsilon}(t)\geq u_{min}\\ 0&\text{ if }u_{min}\leq u_{\upsilon}(t)\leq u_{max}\\ u_{max}-\frac{u_{\upsilon}(t)}{|u_{\upsilon}(t)|+1}&\text{ if }u_{\upsilon}(t)\leq u_{max}\end{cases}\end{aligned}$ (20)
@@ V-B The GRC Framework
$e_{j}(t)=x_{j}-x_{jd},\hskip 5.69046ptz_{j}=\begin{cases}e_{j}-u_{0},&\text{ if }j=2\\ e_{j}&\text{else}\end{cases}$ (21)
@@
$u_{\upsilon}=\begin{cases}-\frac{1}{2}\left(k_{\upsilon}+\epsilon_{\upsilon}\hat{\chi}_{\upsilon}\right)z_{\upsilon+1}-z_{\upsilon},&\text{ if }\upsilon=1\\ -\frac{1}{2}\left(k_{\upsilon}+\epsilon_{\upsilon}\hat{\chi}_{\upsilon}\right)z_{\upsilon+1}&\text{else}\end{cases}$ (22)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
where $\upsilon=2$ or $3$ depending on the number of subsystems, which can be two ($k=0,...,2$: number of subsystems is three) or three ($k=0,...,3$: number of subsystems is four), and:
@@
$\Psi=\min[\Psi_{1},...,\Psi_{4}],\hskip 5.69046pt\tilde{\mu}=\tilde{\mu}_{1}+\tilde{\mu}_{2}+\tilde{\mu}_{3}+\tilde{\mu}_{4}$ (63)
@@
$G_{0}\left(\bar{\tau}_{0}\right):=\left\{\|x_{e}\|\leq\bar{\tau}_{0}:=\sqrt{\frac{\frac{2}{S_{min}}\tilde{\mu}\Psi^{-1}}{1-\overset{*}{Q}}}\right\}$ (67)

### v4

@@ V-A Generic Servo-Driven Actuation System Model
where $\alpha_{1}$$\alpha_{4}$ are unknown positive coefficients. $F_{1}$,…, $F_{4}$ are unknown state-variant modeling and uncertainty terms. $D_{1}$$D_{4}$ are time-variant external disturbances and load effects. By defining $\upsilon=0,1,2,3$, $Sat_{\upsilon}(u_{\upsilon})=s_{1\upsilon}u_{\upsilon}(t)+s_{2\upsilon}$ is a constraining function of the control input $u_{\upsilon}$, which was defined as [24]:
[…]
@@ V-B The GRC Framework
[…]
@@
[…]
@@
$\|\alpha_{1}^{-1}F_{1}\|\hskip 2.84544pt\leq\zeta_{0}\hskip 2.84544ptr_{0}\hskip 2.84544pt,\hskip 5.69046pt\|\alpha_{1}^{-1}{D_{1}}^{*}\|\hskip 2.84544pt\leq{d}_{max1}$ (29)
@@
$\Psi_{1}=\min[k_{0},\hskip 5.69046pt\delta_{0}\gamma_{0}],\hskip 5.69046pt\tilde{\mu}_{1}=\frac{1}{4}\lambda_{0}^{-1}+\frac{1}{2}\delta_{0}{{\chi}^{*}_{0}}^{2}$ (40)
@@
$\Psi_{2}=\min[k_{1},\hskip 5.69046pt\delta_{1}\gamma_{1}],\hskip 5.69046pt\tilde{\mu}_{2}=\frac{1}{4}\lambda_{1}^{-1}+\frac{1}{2}\delta_{1}{{\chi}^{*}_{1}}^{2}$ (50)
@@
$\Psi_{3}=\min[k_{2},\hskip 5.69046pt\delta_{2}\gamma_{2}],\hskip 5.69046pt\tilde{\mu}_{3}=\frac{1}{4}\lambda_{2}^{-1}+\frac{1}{2}\delta_{2}{{\chi}^{*}_{2}}^{2}$ (57)
$\Psi_{4}=\min[k_{3},\hskip 5.69046pt\delta_{3}\gamma_{3}],\hskip 5.69046pt\tilde{\mu}_{4}=\frac{1}{4}\lambda_{3}^{-1}+\frac{1}{2}\delta_{3}{{\chi}^{*}_{3}}^{2}$
@@
where $\upsilon=2$ or $3$ depending on the number of subsystems, which can be two (k=0,,2k=0,...,2: number of subsystems is three) or three (k=0,,3k=0,...,3: number of subsystems is four), and:
@@
$\Psi=\min[\Psi_{1},.,\Psi_{4}],\hskip 5.69046pt\tilde{\mu}=\tilde{\mu}_{1}+\tilde{\mu}_{2}+\tilde{\mu}_{3}+\tilde{\mu}_{4}$ (63)
@@
$G_{0}\left(\bar{\tau}_{0}\right)=\left\{\|x_{e}\|\leq\bar{\tau}_{0}:=\sqrt{\frac{\frac{2}{S_{min}}\tilde{\mu}\Psi^{-1}}{1-\overset{*}{Q}}}\right\}$ (67)